OpenStudy (anonymous):

A neutron collides elastically with a helium nucleus (at rest initially) whose mass is four times that of the neutron. The helium nucleus is observed to rebound at an angle '2 = 43° from the neutron's initial direction. The neutron's initial speed is 5.4 105 m/s. Determine the angle at which the neutron rebounds, '1, measured from its initial direction. ° What is the speed of the neutron after the collision? m/s What is the speed of the helium nucleus after the collision?

OpenStudy (anonymous):

Conservation of momentum means that the initial momentum of the neutron will be split between it and the helium atom after the collision. Call that initial direction the x direction, then after collision all that momentum goes into the x-direction components of momentum for the neutron and the helium atom. The initial y-momentum was zero, so the sum of the post-collision components of y momentum will be zero. Elastic means that the kinetic energy is conserved, so that the (1/2)mv^2 of the neutron initially will be shared by the neutron and the helium atom afterward.

OpenStudy (anonymous):

How would I be able to find theta though? And since it is elastic the neutron and helium would have the same speed?

OpenStudy (anonymous):

Elastic only means energy is conserved, not that they have the same speed. You get theta by looking at the components of velocity in the x and the y direction. Rebounding at a velocity v at a theta=43o angle from the x direction would mean helium having an x component of velocity v cos (43) and a y component of v sin(43). Try to work it out and I will spend a little time on it, too. Make life easier by setting m1 of neutron = 1, m2 of helium=4. Their ratios should be what is important in this problem, not their absolute masses..

OpenStudy (anonymous):

Well, it gets messy. You have three equations: two for momentum (x, y directions) and one for energy. You know the masses, and the initial velocity, and one final angle (43o), but you have three unknowns: theta and the resulting two velocities. I have not been able to boil them down yet. x momentum: m v1 = M v2 cos(43) + m v1' cos(theta) y momentum: 0 = M v2 sin(43) +m v1' sin(theta) energy: (1/2)m v1^2 = (1/2) M v2^2 + (1/2) m v1'^2 Using M/m=4 and rearranging, the energy equation gives v1^2 - v1'^2 = 4 v2^2, the energy lost by the neutron goes to He. The y momentum equation shows that sin(theta) = - M v2 sin(43) / m v1' = (-4 v2 / v1') sin(43) so that theta is directed somewhat below the horizontal. Brute force with the momentum equations will let you get rid of the terms with M v2 in them and get v1' and theta, probably, but I am stopping here!

OpenStudy (anonymous):

Thanks for all the help!